# MIRROR: A Novel Approach for the Automated Evaluation of Open-Ended Question Generation

### v1

@@ 1 Introduction
Automated question generation (AQG) is crucial in education because it enhances critical thinking, promotes active learning, and provides personalized learning experiences. Currently, metrics such as BLEU, METEOR, and ROUGE are used for the purpose of evaluation of open ended questions (Zhang et al., 2021; Al Faraby et al., 2023; Deroy et al., 2023a, 2024c). However, traditional automated evaluation metrics such as BLEU, ROUGE, or METEOR, which are often used for tasks like machine translation or summarization, may not be well-suited for evaluating the quality of generated questions (Alva-Manchego et al., 2021; Bhandari et al., 2020; Reiter, 2018; Rao et al., 2022; Deroy et al., 2021). These metrics primarily measure surface-level similarity to reference questions, rather than deeper aspects of question quality such as relevance, clarity, and engagement (Nema and Khapra, 2018; Maity et al., 2024e). Some aspects of question quality, such as engagement, pedagogical value, and the potential to stimulate critical thinking, require a human-like understanding and judgment that automated systems cannot yet replicate. However human judgements are costly and cannot be replicated over large samples of generated questions. So we propose an LLM feedback-based system called MIRROR (Multi-LLM Iterative Review and Response for Optimized Rating) for automating the process of human evaluation for questions generated by AQG systems. We understand that LLMs have understanding and reasoning capabilities.
@@ 2 Related Work
Evaluating open-ended question generation using automated metrics (e.g., BLEU, ROUGE, METEOR, etc.) presents significant challenges (Nema and Khapra, 2018; Gong et al., 2022; Ushio et al., 2022; Maity et al., 2024a, 2023, f, g; Deroy and Maity, 2023b; Nigam et al., 2023). These metrics often fail to capture higher-order cognitive skills and overlook deeper educational values such as stimulating critical thinking (Al Faraby et al., 2023; Mulla and Gharpure, 2023; Deroy et al., 2024b; Nigam and Deroy, 2023; Deroy et al., 2023b; Maity et al., 2024c; Deroy and Maity, 2023a; Deroy et al., 2024a). Automated evaluations may not fully account for nuances such as context relevance and cognitive complexity, making it difficult to assess whether questions promote skills such as analysis, synthesis, and evaluation (Mulla and Gharpure, 2023; Deroy and Maity, 2024a, b, c, d; Ghosh et al., 2023; Maity and Deroy, 2023). Furthermore, scalability issues arise because human evaluation for large datasets is impractical (Al Faraby et al., 2023; Deroy, 2019; Maity et al., 2024d). Using LLMs to evaluate human-like criteria such as grammaticality, relevance, appropriateness, novelty, and complexity has been a prominent research area (Chiang and Lee, 2023). Aligning machine-generated content with human judgment is crucial for practical applicability. Incorporating feedback loops within LLMs to refine their outputs is an emerging field (Madaan et al., 2023; Chaudhari et al., 2024; Ouyang et al., 2022; Stiennon et al., 2020). Techniques like Reinforcement Learning from Human Feedback (RLHF) fine-tune LLMs based on human preferences, enhancing the quality and contextual relevance of responses (Yuan et al., 2023; Chaudhari et al., 2024; Wu et al., 2023). This work proposes a novel method using LLM feedback to evaluate open-ended questions based on human-like metrics: grammaticality (Ushio et al., 2022; Maity et al., 2024b, a), relevance (Maity et al., 2024b), appropriateness (Maity et al., 2024b), complexity (Gong et al., 2022; Maity et al., 2024b), and novelty (Maity et al., 2024b). This approach aims to bridge the gap between automated metrics and human evaluation, enhancing the quality and applicability of generated questions in educational settings.
@@ 4 Methodology
In this section, we discuss the generation of questions via prompting GPT-3.5 Turbo (Section 4.1), the direct approach to evaluating the quality of the generated questions (Section 4.2), the feedback-based approach for evaluating generated question quality (Section 4.3), and the correlation between the best-performing LLM and human experts (Section 4.4).

### v3

@@ 1 Introduction
Automated question generation (AQG) is crucial in education because it enhances critical thinking, promotes active learning, and provides personalized learning experiences. Currently, metrics such as BLEU, METEOR, and ROUGE are used for the purpose of evaluation of open ended questions (Zhang et al., 2021; Al Faraby et al., 2023). However, traditional automated evaluation metrics such as BLEU, ROUGE, or METEOR, which are often used for tasks like machine translation or summarization, may not be well-suited for evaluating the quality of generated questions (Alva-Manchego et al., 2021; Bhandari et al., 2020; Reiter, 2018). These metrics primarily measure surface-level similarity to reference questions, rather than deeper aspects of question quality such as relevance, clarity, and engagement (Nema and Khapra, 2018). Some aspects of question quality, such as engagement, pedagogical value, and the potential to stimulate critical thinking, require a human-like understanding and judgment that automated systems cannot yet replicate. However human judgements are costly and cannot be replicated over large samples of generated questions. So we propose an LLM feedback-based system called MIRROR (Multi-LLM Iterative Review and Response for Optimized Rating) for automating the process of human evaluation for questions generated by AQG systems. We understand that LLMs have understanding and reasoning capabilities.
@@ 2 Related Work
Evaluating open-ended question generation using automated metrics (e.g., BLEU, ROUGE, METEOR, etc.) presents significant challenges (Nema and Khapra, 2018; Gong et al., 2022; Ushio et al., 2022; Maity et al., 2024a). These metrics often fail to capture higher-order cognitive skills and overlook deeper educational values such as stimulating critical thinking (Al Faraby et al., 2023; Mulla and Gharpure, 2023). Automated evaluations may not fully account for nuances such as context relevance and cognitive complexity, making it difficult to assess whether questions promote skills such as analysis, synthesis, and evaluation (Mulla and Gharpure, 2023). Furthermore, scalability issues arise because human evaluation for large datasets is impractical (Al Faraby et al., 2023). Using LLMs to evaluate human-like criteria such as grammaticality, relevance, appropriateness, novelty, and complexity has been a prominent research area (Chiang and Lee, 2023). Aligning machine-generated content with human judgment is crucial for practical applicability. Incorporating feedback loops within LLMs to refine their outputs is an emerging field (Madaan et al., 2023; Chaudhari et al., 2024; Ouyang et al., 2022; Stiennon et al., 2020). Techniques like Reinforcement Learning from Human Feedback (RLHF) fine-tune LLMs based on human preferences, enhancing the quality and contextual relevance of responses (Yuan et al., 2023; Chaudhari et al., 2024; Wu et al., 2023). This work proposes a novel method using LLM feedback to evaluate open-ended questions based on human-like metrics: grammaticality (Ushio et al., 2022; Maity et al., 2024b, a), relevance (Maity et al., 2024b), appropriateness (Maity et al., 2024b), complexity (Gong et al., 2022; Maity et al., 2024b), and novelty (Maity et al., 2024b). This approach aims to bridge the gap between automated metrics and human evaluation, enhancing the quality and applicability of generated questions in educational settings.
@@ 4 Methodology
In this section, we discuss the generation of questions via prompting GPT-3.5 Turbo (Section 4.1), the direct approach to evaluating the quality of the generated questions (Section 4.2), the feedback-based approach for evaluating generated question quality (Section 4.3), and the correlation between the best-performing LLM and human experts (Section 4.4). In our experiments, we used state-of-the-art LLMs, such as GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and Llama2-70b (Touvron et al., 2023), to assess both direct and feedback-based approaches.